If R and H are the horizontal range and maximum height attained by a projectile, then its speed of projection ??

Grade:12th pass

1 Answers

Vikas TU
7 years ago
Range R = 2u/g
and
Max. Height, H =u^2/2g
Divide the Range and Height we get,
R/H = 4/u
u = 4H/R m/s is the speed of projection.